A ball is dropped from a hot air ballon moving vertically upward with a speed of30m/s. If the balloon is 80m above the ground at the instant the ball is dropped. How long dose it take for the ball to reach the ground?

Grade:12th pass

1 Answers

Akash Biswas
6 years ago
Firstly we have, t1=u/g or t1 = 30/10=3s H=u^2 /2g or H = 30*30/10=45Secondly,Total distance H=45+80=125T=(2h/g)^1/2T=5s Total time =t1+T=8s